OpenStudy (alana242):

The area of the centre circle on a football pitch is 265 m(squared). Calculate the radius of the circle to the nearest 0.1 m. Someone please help

OpenStudy (calculusxy):

\[Area = \pi r^2 \] \[265 = \pi r^2 \] \(\pi\) can be 3.14 \[265 = 3.14 r^2 \] divide both sides by 3.14 and then find the square root of the quotient to find the radius

OpenStudy (alana242):

I got 9.2? @calculusxy

OpenStudy (alana242):

@calculusxy

OpenStudy (calculusxy):

correct!